Hi! My Algebra teacher has not been here the past two weeks so I have been teaching myself everything. I can't seen to figure out how to solve equations by completing the square. For example: f^2+3f=88 Thank you!
Answer by scott8148(6628)

if the coefficient of the squared term is one (as in this case)

the constant term for the "perfect" square is ___ half of the 1st order coefficient, squared

f^2 + 3f + (3/2)^2 = 88 + (3/2)^2

(f + 3/2)^2 = 88 + (9/4) = 361/4

taking square root ___ f + 3/2 = ħsqrt(361/4) ___ (don't forget ħ)

f + 3/2 = ħ19/2

f = 19/2 - 3/2 = 8 , f = -19/2 - 3/2 = -11


if the coefficient of the squared term is not one; then the 1st step is to divide by the coefficient